A firm has determined its cost of each source of capital and optimal capital structure which is composed of the following sources and target market value proportions.

Source of capital Target Market Proportions After-tax Cost
Long-term Debt 35% 9%
Preferred Stock 10 14
Common Stock Equity 55 20


The firm is considering an investment opportunity, which has an internal rate of return of 18 percent. The project

should not be considered because its internal rate of return is less than the cost of long-term debt.

should be considered because its internal rate of return is greater than the cost of debt.

should not be considered because its internal rate of return is less than the weighted average cost of capital.

should be considered because its internal rate of return is greater than the weighted average cost of capital.
Answer #1

Se WACC ( cost of capital) ko We x kg) + (Wd xkd) + (Wpx Kp) | = ( 5SX02) +(0, 35 K001)+(0.70 10.1 % = 0.11 +0.0315 + 0.014 =